# Adjusting for Heavy Censoring and Double-Dipping to Compare Risk Stratification Abilities of Existing Models for Time to Diagnosis of Huntington Disease

### v1

@@ 3.1 Data Description
To compare the four time-to-diagnosis models, we turned to data from the ENROLL-HD study, which were not used to develop any of the models. ENROLL-HD is a longitudinal observational study developed to improve scientific understanding of HD. Patients from three risk groups were enrolled in this study: (i) those who were clinically diagnosed with HD, (ii) those with CAG repeats $\geq 36$ but who were not yet diagnosed, and (iii) those with CAG repeats $\leq 36$ Sathe2021. The study is massive in scope, with $21818$ current patients enrolled across $156$ clinical sites (mostly in Europe and North America, with some in Latin America and Oceania). Given its broad patient population and global scope, ENROLL-HD serves as a powerful research platform for studying HD.
@@ 3.2 Comparing Risk Stratification Abilities
We present two comparative analyses of the four models. First, we evaluated them using the published distributions (if applicable) and covariates but re-estimated parameters based on ENROLL-HD. Second, we applied them with the original published parameters. The first comparison assessed the generalizability of the models’ structure to new data and the second gauged their performance with the original estimates based on other studies.
@@ 3.2.1 Uno’s C Statistics
To evaluate the performance with newly-estimated parameters tailored to the ENROLL-HD data, we used a 5-fold cross-validation procedure. We randomly split the data into $5$ mutually exclusive subgroups (“folds”). Selection of the folds was stratified so that each fold contained $725/5=145$ uncensored patients, preserving the sample’s censoring rate ($78\%$) within each.
@@
All four models demonstrated moderate-to-strong risk stratification abilities overall (Figure 1), with global Uno’s C statistics narrowly ranging from $0.80$ (CAP model) to $0.86$ (MRS model). Interestingly, the models’ ordering was very consistent across all time points. The MRS model outperformed PIN, which outperformed the nearly-tied CAP and Langbehn models. We also saw that all models tended to perform worse over time, with each model’s maximum Uno’s C statistic observed at $\tau=1$ year from enrollment. This gradual decline in performance could be because the covariates were only taken at baseline (i.e., treated as time-independent). For example, the PIN model does not update predictions as more recent SDMT scores are collected, even though those values would be more informative over time.
Figure 1: Time-specific Uno’s C statistics by model (estimated parameters) and evaluation time, with statistic given by mean from 5-fold cross-validation. The global Uno’s C statistic is a weighted average of the time-specific values. We considered four time-to-diagnosis models: CAG-Age Product (CAP), Langbehn, Multivariate Risk Score (MRS), and Prognostic Index Normed (PIN).
@@
In addition to comparing these models using 5-fold cross-validation, we evaluated them with the published parameters applied to the entire analytic sample (Web Appendix A). We expected there to be differences in performance between the cross-validation and published parameter analyses, as the published parameters were obtained using different datasets. However, the differences in performance between the two analyses were within 1% to 6%, suggesting that the published parameters are externally valid.
@@ 3.2.2 Kaplan-Meier Adjusted ROC Curves
Uno’s C statistics assessed how well the models rank pairs of patients in terms of log-risk scores. We also compared the models in terms of the AUC for their Kaplan-Meier-adjusted ROC curves. While both metrics assess risk stratification, AUC differs from Uno’s C statistic in that it provides an assessment of how well the models separate the group of patients diagnosed by time $\tau$ from the group undiagnosed by time $\tau$, whereas Uno’s C statistic describes how well pairs of patients are ranked. As in the prior section, we calculated AUC values using 5-fold cross validation to assess the models’ performance when applied to data external to the training set.
The models all produced AUC values similar to their Uno’s C statistics, indicating similarly strong group-based risk stratification (Figure 2). The ordering of models based on their AUC values also matched the ordering based on Uno’s C statistics, with MRS performing best, PIN performing second best, and Langbehn and CAP nearly tied.
Figure 2: Time-specific area under the curve (AUC) of the receiver operating characteristic curve by model (estimated parameters) and evaluation time, with statistic given by mean from 5-fold cross-validation. We considered four time-to-diagnosis models: Langbehn, CAG-Age Product (CAP), Multivariate Risk Score (MRS), and Prognostic Index Normed (PIN).
@@ 3.3 Selecting an Optimal Log-Risk Score Threshold for Clinical Trial Sample Enrichment
We leveraged Kaplan-Meier-adjusted ROC curves to inform sample enrichment for preventative clinical trials of HD. We used ROC curves to select an optimal log-risk score threshold, a necessary step toward sample enrichment, because they allowed us to consider the predictive accuracy of different possible thresholds (unlike Uno’s C statistic). For simplicity and to avoid double-dipping, we used the published parameters to demonstrate this procedure. If, instead, parameters were to be re-estimated, cross-validation would be needed here, as well. (Cross-validation could result in different thresholds and recommended sample sizes in each fold, which would need to be combined for implementation.)
We exclude the Langbehn model from this portion of our analysis because it was difficult to implement and was tied with the CAP in terms of risk stratification. We considered four hypothetical trial durations ranging from 2255 years. For each combination of model and trial duration, we first calculated the optimal log-risk score threshold. Using the Kaplan-Meier-adjusted ROC curve, we identified an optimal threshold $q^{*}$ that maximizes Youden’s J statistic at time $t$, given by $J(q,t)=[\widehat{\textrm{TPR}}(q,t)+\{1-\widehat{\textrm{FPR}}(q,t)\}-1]$ Youden1950. The enriched trial would enroll only patients whose estimated log-risk scores exceed $q^{*}$.
Then, for the optimal threshold, we estimated the expected diagnosis rate in the untreated arm (i.e., the proportion of patients who would be diagnosed with HD during the trial) among those whose who would be enrolled. With this expected rate, we calculated the per-arm sample sizes required to detect treatment effects of $30\%$, $40\%$, and $50\%$ reductions in the rate of HD diagnosis, assuming $80\%$ power and a two-sided type-I error rate of $\alpha=0.05$. This sample enrichment strategy follows the same framework as Paulsen et al. (2019) Paulsen2019.
@@
The optimal threshold for a $3$-year clinical trial was estimated to be $8.96$ for the MRS model. Therefore, we would aim to recruit patients whose estimated log-risk scores were $\geq 8.96$. In doing so, we would expect $42\%$ of the untreated group to be diagnosed with HD during the trial. We estimated that $77$ patients per arm would be required to detect a 50% treatment effect with $80\%$ power and $\alpha=0.05$. For smaller treatment effects of $40\%$ and $30\%$, we obtained per-arm sample size requirements of $124$ and $227$ patients, respectively.
In prior work, Paulsen et al. (2019) Paulsen2019 estimated smaller per-arm sample sizes for each of these effect sizes: $67$ at $50\%$, $103$ at $40\%$, and $179$ at $30\%$. Although we used the same published parameter estimates for the MRS model, our analysis differs from that of Paulsen et al. (2019) Paulsen2019 in two key ways. First, we used data from ENROLL-HD, whereas Paulsen et al. (2019) Paulsen2019 used PREDICT-HD, the same dataset on which the CAP, MRS, and PIN models were originally developed to some extent. Second, our ROC analysis adjusted for censoring, which was crucial given that the ENROLL-HD data were $78\%$ censored, whereas the prior work did not. Compared to our censoring-adjusted, externally-validated calculations, the sample sizes estimates from Paulsen et al. (2019) Paulsen2019 would result in an underpowered clinical trial, thereby increasing the probability of failing to detect a truly effective HD treatment.
@@ 4 Discussion
In comparing model performance via 5-fold cross-validation (Section 3.2), we observed a clear trade-off between risk stratification and parsimony. The MRS model demonstrated the strongest risk stratification capability but was the most complex (requiring the most covariates [eight]). Meanwhile, the PIN model, which required half as many covariates (four), performed only $3\%$ worse than MRS. The CAP and Langbehn models, using only two covariates, tie for third place, each performing about $4\%$ worse than PIN. However, based on our experiences, we recommend against using the Langbehn model: It is the most computationally complex and is no more parsimonious than the CAP model since both require the same covariates yet their performance is nearly identical. If parsimony is a priority, the CAP model is simpler to implement and interpret. If more clinical measurements are available, the MRS and PIN models outperform both. Crucially, our ordering matched that of Paulsen et al. (2019) Paulsen2019, so our work strengthens their arguments.
@@
Estimating the CAP, MRS, and PIN models was straightforward using the survival package in R Therneau2022. Estimating the Langbehn model was more difficult because of its unique structure, and no off-the-shelf software existed. We implemented an MLE procedure for this model in R, which is available online at www.github.com/sarahlotspeich/hd_ttd_modeling. A grid search over possible values of one of the parameters ($\beta_{2}$ in (1)) was needed, since it would otherwise converge to its lower bound during optimization.
Since we focus on sample enrichment for preventative clinical trials, where we recruit patients that are nearer to diagnosis than others, the models’ ability to stratify potential patients by risk was paramount. However, one could also compare these models in terms of their calibration, which refers to the strength of agreement between a model’s predicted survival probabilities and the observed rates of diagnosis. Park et al. (2021) Park2021 provide a thorough review of metrics for assessing a survival model’s calibration. Calibration plots provide an intuitive, graphical assessment. The Brier score, which simultaneously quantifies both risk stratification and calibration, could also be used, though it would have to be adjusted for censoring (e.g., through IPCW) if applied to HD data.
When calculating the optimal thresholds and per-arm sample sizes for a preventative clinical trial, we applied each model’s published parameter estimates to our entire analytic sample. Performing these calculations with a cross-validation procedure, similar to Section 3.2, would allow us to estimate the variability of the estimated sample sizes.
Another avenue for future development is to use a survival model with time-dependent covariates (e.g., changing TMS and SDMT scores over time). As we discussed when comparing the models’ performance via cross-validation (Section 3.2), they tended to perform worse at later time points. This later-stage performance could potentially be improved by incorporating time-dependent covariates into the models, such as updating a patient’s estimated log-risk score when new TMS values are available.

### v2

@@ 3.1 Data Description
To compare the four time-to-diagnosis models, we turned to data from the ENROLL-HD study, which were not used to develop any of the models. ENROLL-HD is a longitudinal observational study developed to improve scientific understanding of HD. Patients from three risk groups were enrolled in this study: (i) those who were clinically diagnosed with HD, (ii) those with CAG repeats $\geq 36$ but who were not yet diagnosed, and (iii) those with CAG repeats $\leq 36$ (Sathe et al., 2021). The study is massive in scope, with $21818$ current patients enrolled across $156$ clinical sites (mostly in Europe and North America, with some in Latin America and Oceania). Given its broad patient population and global scope, ENROLL-HD serves as a powerful research platform for studying HD.
@@ 3.2 Comparing Risk Stratification Abilities
We present two comparative analyses of the four models. First, we applied each model with the original published parameters (external validation). Second, we evaluated each using the published distributions (if applicable) and covariates but updated parameters based on ENROLL-HD (model updating). The first comparison gauged the models’ transportability to new HD studies, and the second assessed whether tailoring coefficients to new data but retaining the original structure could further improve performance. In each of the following subsections, we begin by discussing cross-validation with the updated models and end by comparing them to the originals. Ultimately, we found that all models performed similarly with original or updated coefficients.
@@ 3.2.1 Uno’s C Statistics
To evaluate the performance with updated parameters tailored to the ENROLL-HD data, we used a $5$-fold cross-validation procedure. We randomly split the data into $5$ mutually exclusive subgroups (“folds”). Selection of the folds was stratified so that each fold contained $725/5=145$ uncensored patients, preserving the sample’s censoring rate ($78\%$) within each.
@@
All four models demonstrated moderate-to-strong risk stratification abilities overall (Figure 1(B)), with global Uno’s C statistics narrowly ranging from $0.80$ (CAP model) to $0.86$ (MRS model). Interestingly, the models’ ordering was very consistent across all time points. The MRS model outperformed PIN, which outperformed the nearly-tied CAP and Langbehn models. We also saw that all models tended to perform worse over time, with each model’s maximum Uno’s C statistic observed at $\tau=1$ year from enrollment. This gradual decline in performance could be because the covariates were only taken at baseline (i.e., treated as time-independent). For example, the PIN model does not update predictions as more recent SDMT scores are collected, even though those values would be more informative over time.
Figure 1: Time-specific Uno’s C statistics by model and evaluation time for both published parameters (A) and updated parameters (B), with statistic given by mean from 55-fold cross-validation for updated parameters. The average Uno’s C statistic is a weighted average of the time-specific values. We considered four time-to-diagnosis models: CAG-Age Product (CAP), Langbehn, Multivariate Risk Score (MRS), and Prognostic Index Normed (PIN).
@@
We also evaluated the four models with the published parameters applied to the entire analytic sample (Figure 1(A)). We expected there to be differences in performance between the updated and published parameter analyses, as the published parameters were obtained using different datasets. However, the differences in Uno’s C between the two analyses were within $1\%$ to $6\%$ across all time points, suggesting that the updated parameters did not noticeably improve model performance.
@@ 3.2.2 Kaplan-Meier Adjusted ROC Curves
As in the prior section, we calculated AUC values using $5$-fold cross validation to assess the models’ performance when applied to data external to the training set. The models all produced AUC values similar to their Uno’s C statistics, indicating similarly strong group-based risk stratification (Figure 2). The ordering of models based on their AUC values also matched the ordering based on Uno’s C statistics, with MRS performing best, PIN performing second best, and Langbehn and CAP nearly tied.
Figure 2: Time-specific area under the curve (AUC) of the receiver operating characteristic curve by model and evaluation time for both published parameters (A) and updated parameters (B). We considered four time-to-diagnosis models: Langbehn, CAG-Age Product (CAP), Multivariate Risk Score (MRS), and Prognostic Index Normed (PIN).
@@ 3.3 Selecting an Optimal Log-Risk Score Threshold for Clinical Trial Sample Enrichment
We leveraged Kaplan-Meier-adjusted ROC curves to inform sample enrichment for preventative clinical trials of HD. We used ROC curves to select an optimal log-risk score threshold, a necessary step toward sample enrichment, because they allowed us to consider the predictive accuracy of different possible thresholds (unlike Uno’s C statistic). For simplicity, we used the published parameters to demonstrate this procedure. If, instead, parameters were to be updated, cross-validation would be needed here, as well. (Cross-validation could result in different thresholds and recommended sample sizes in each fold, which would need to be combined for implementation.)
We exclude the Langbehn model from this portion of our analysis because it was difficult to implement and was tied with the CAP in terms of risk stratification. We considered four hypothetical trial durations ranging from 2255 years. For each combination of model and trial duration, we first calculated the optimal log-risk score threshold. Using the Kaplan-Meier-adjusted ROC curve, we identified an optimal threshold $q^{*}$ that maximizes Youden’s J statistic at time $t$, given by $J(q,t)=[\widehat{\textrm{TPR}}(q,t)+\{1-\widehat{\textrm{FPR}}(q,t)\}-1]$ (Youden, 1950). The enriched trial would enroll only patients whose estimated log-risk scores exceed $q^{*}$.
Then, for the optimal threshold, we estimated the expected diagnosis rate in the untreated arm (i.e., the proportion of patients who would be diagnosed with HD during the trial) among those whose who would be enrolled. With this expected rate, we calculated the per-arm sample sizes required to detect treatment effects of $30\%$, $40\%$, and $50\%$ reductions in the rate of HD diagnosis, assuming $80\%$ power and a two-sided type-I error rate of $\alpha=0.05$. This sample enrichment strategy follows the same framework as Paulsen et al. (2019) Paulsen et al. (2019).
@@
The optimal threshold for a $3$-year clinical trial was estimated to be $8.96$ for the MRS model. Therefore, we would aim to recruit patients whose estimated log-risk scores were $\geq 8.96$. In doing so, we would expect $42\%$ of the untreated group to be diagnosed with HD during the trial. We estimated that $77$ patients per arm would be required to detect a $50\%$ treatment effect with $80\%$ power and $\alpha=0.05$. For smaller treatment effects of $40\%$ and $30\%$, we obtained per-arm sample size requirements of $124$ and $227$ patients, respectively.
In prior work, Paulsen et al. (2019) Paulsen et al. (2019) estimated smaller per-arm sample sizes for each of these effect sizes: $67$ at $50\%$, $103$ at $40\%$, and $179$ at $30\%$. Although we used the same published parameter estimates for the MRS model, our analysis differs from that of Paulsen et al. (2019) Paulsen et al. (2019) in two key ways. First, we used data from ENROLL-HD, whereas Paulsen et al. (2019) Paulsen et al. (2019) used PREDICT-HD, the same dataset on which the CAP, MRS, and PIN models were originally developed to some extent. Second, our ROC analysis adjusted for censoring, which was crucial given that the ENROLL-HD data were $78\%$ censored, whereas the prior work did not. Compared to our censoring-adjusted, externally-validated calculations, the sample sizes estimates from Paulsen et al. (2019) Paulsen et al. (2019) would result in an underpowered clinical trial, thereby increasing the probability of failing to detect a truly effective HD treatment.
@@ 4 Discussion
In comparing model performance via $5$-fold cross-validation (Section 3.2), we observed a clear trade-off between risk stratification and parsimony. The MRS model demonstrated the strongest risk stratification capability but was the most complex (requiring the most covariates [eight]). Meanwhile, the PIN model, which required half as many covariates (four), performed only $3\%$ worse than MRS. The CAP and Langbehn models, using only two covariates, tie for third place, each performing about $4\%$ worse than PIN. However, based on our experiences, we recommend against using the Langbehn model: It is the most computationally complex and is no more parsimonious than the CAP model since both require the same covariates yet their performance is nearly identical. If parsimony is a priority, the CAP model is simpler to implement and interpret. If more clinical measurements are available, the MRS and PIN models outperform both. Crucially, our ordering matched that of Paulsen et al. (2019) Paulsen et al. (2019), so our work strengthens their arguments.
@@
Estimating the CAP, MRS, and PIN models was straightforward using the survival package in R (Therneau, 2022). Estimating the Langbehn model was more difficult because of its unique structure, and no off-the-shelf software existed. We implemented an MLE procedure for this model in R, which is available online at www.github.com/sarahlotspeich/hd_ttd_modeling. A grid search over possible values of one of the parameters ($\beta_{2}$ in (1)) was needed, since it would otherwise converge to its lower bound during optimization.
Since we focus on sample enrichment for preventative clinical trials, where we recruit patients that are nearer to diagnosis than others, the models’ ability to stratify potential patients by risk was paramount. However, one could also compare these models in terms of their calibration, which refers to the strength of agreement between a model’s predicted survival probabilities and the observed rates of diagnosis. Park et al. (2021) Park et al. (2021) provide a thorough review of metrics for assessing a survival model’s calibration. Calibration plots provide an intuitive, graphical assessment. The Brier score, which simultaneously quantifies both risk stratification and calibration, could also be used, though it would have to be adjusted for censoring (e.g., through IPCW) if applied to HD data.
When calculating the optimal thresholds and per-arm sample sizes for a preventative clinical trial, we applied each model’s published parameter estimates to our entire analytic sample. Performing these calculations with the updated parameters via a cross-validation procedure, similar to Section 3.2, would also be straightforward.
Another avenue for future development is to build upon the best-performing model identified here. For example, all models tended to perform worse at later time points. This later-stage performance could potentially be improved by incorporating time-dependent covariates into the models (e.g., changing TMS and SDMT scores over time). Further, in preventative clinical trials, patients are recruited pre-diagnosis, which largely eliminates concerns about competing risks. However, trials targeting patients at later disease stages would need to consider incorporating a competing risk analysis (e.g., for censoring due to death from HD-related complications before receiving a formal clinical diagnosis).